OpenStudy (anonymous):

Solve: x-5=√x+7 a) x = 2 or x = 9 b) x = 9 c) No Solution d) x = 2 The square root is over the x + 7

OpenStudy (anonymous):

b) x = 9

OpenStudy (anonymous):

square both sides (x-5)^2=x+7 x^2-10x+25=x+7 x^2-11+18=0 (x+2)(x-9)=0 x=-2, x=9 x=9